# itemizestditemize

### v1

@@ IV-B Vision-based Planar Pushing in Simulation
Policy. We train a vision-based diffusion policy ${\cal P}(\cdot)$ [15] using ResNet18 as the vision encoder and UNet as the action diffusion network. We used $500$ expert demonstrations to train the policy for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon of 2, prediction horizon of 16, and action horizon of 9.
@@
World model. As in §III-B, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. Same as in §IV-A, the training data includes random explorations.
@@
Reward. Different from the state-based planar pushing, since the future predictions are images, we cannot directly compute a reward signal. Therefore, we first train a pose prediction network with ResNet18 plus MLP, and then compute the registration loss using the predicted object pose. Note that this learned reward model ${\cal R}(\cdot)$ is still differentiable.
@@
Online planning: With the vision-based policy, the visual world model, and the learned reward model, we follow the same online planning protocol described in §IV-A.
Quality of visual world modeling. Before presenting the results of GPC for online planning, we first study whether our visual world modeling technique is better than some baselines. Particularly, we chose two baselines, one from the pioneering deep visual foresight [20], and the other from the more recent AVDC paper [19]. The world model in deep visual foresight leverages CNN and LSTM for future prediction, and we follow the Github implementation. [^3] https://github.com/Xiaohui9607/physical_interaction_video_prediction_pytorch AVDC is originally designed for video diffusion conditioned on language, and we modify its official implementation [^4] https://github.com/flow-diffusion/AVDC to condition on robot actions. Notably, AVDC is a multiple-step world model as it simultaneously diffuses future images together, unlike our recursive structure. Fig. 4 compares the quality of future predictions of three methods (ours, deep visual foresight, and AVDC) against the groundtruth in two examples. We see that diffusion-based visual world modeling (ours and AVDC) clearly dominates the classical deep visual foresight. Between ours and AVDC, we can observe that the future predictions from our method are better aligned with the groundtruth.
Planning performance. Same as §IV-A, we study the impact of $K$ and $M$ on the performance of GPC, its comparison with Offline RL, and the importance of random exploration.
In Table III, we study how different variations of $K$ and $M$ affect the performance of GPC. We can see that (a) GPC-RANK increases around $15\%$ from the diffusion policy baseline (when $K=1,M=0$); (b) GPC-OPT increases around $25\%$ from the diffusion policy baseline; and (c) GPC-RANK+OPT increases up to $40\%$ from the baseline. Contrasting the vision-based results with the state-based results, we observe that GPC-OPT appears to have a larger advantage in vision-based push-T. This indeed makes sense because the baseline vision-based diffusion policy only has a score of $0.642$, leaving a lot of room for optimization to find better actions.
In the first column of Table III, we show the results of Offline RL. Unfortunately, Offline RL still does not perform as well as even the baseline vision-based diffusion policy.
In Table IV, we compare the performance of GPC-RANK and GPC-OPT when trained with or without random exploration. Consistent with our observations in IV-A, without training on random exploration, the performance of online planning significantly decreases.
Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT $K=100$ $K=1,M=0$ $K=50,M=0$ $K=100,M=0$ $K=200,M=0$ $K=1,M=25$ $K=5,M=10$ $K=10,M=25$ Score $0.620$ $0.642$ $0.698$ $0.739$ $0.719$ $0.791$ 0.824 0.882
TABLE III: Vision-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over $100$ evaluation seeds. $K$ is the number of action proposals and $M$ is the number of gradient steps.
GPC-RANK GPC-OPT $K=100,M=0$ $K=1,M=25$ Score (using world model trained without random exploration) $0.647$ 0.652 Score (using world model trained with random exploration) $0.739$ $0.791$
TABLE IV: The importance of random exploration in world model learning (vision-based push-T).
IV-C Real-world Vision-based Planar Pushing
We then test GPC in real-world planar pushing. We not only show GPC is effective in “plain push-T”, but also show it can work for push-T with collision objects.
Figure 5: Real-world setup.
Setup. Fig. 5 depicts our real-world setup for vision-based planar pushing. The pusher is attached to the end-effector of a Franka Panda robotic arm to push a T block into the target position. An overhead Intel RealSense camera is used to provide image observations to the pusher. To train the reward model, we need to obtain block information as training data so we attached three AprilTags [46] to the T-block, which enables accurate pose estimation from image frames even in the case of occlusion. To collect real-world expert demonstrations, we follow Chi et al. [15] and use a SpaceMouse to teleoperate the robotic arm to push the T block. A and R blocks are the objects that T may collide into during pushing.
IV-C1 Plain push-T
We first test on plain push-T (i.e., without collision with other objects). We train the policy with $100$ expert push-T demonstrations using the architecture mentioned in §IV-B. Because of the slower moving speed of real-world pusher, we use a larger prediction horizon $T=32$. Same as §IV-B, we train the world model using all expert demonstrations and a few random explorations. For the reward model, we rely on AprilTag [46] to obtain the ground-truth object poses and use them to train the reward model ${\cal R}(\cdot)$. To be clear, our GPC pipeline in real world is still purely vision-based and the AprilTags are only used to obtain training data for the reward model (training the pose predictor). We follow the same online planning protocol described in §IV-A. We test GPC-RANK and GPC-OPT on a set of 10 push-T tasks where the T block is initialized at positions not seen during training.
Performance. While the baseline model succeeds 5 out of 10 times, GPC-RANK succeeds 7 out of 10 times and GPC-OPT succeeds 7 out of 10 times. Fig. 10(a) shows the trajectories for one test. More testing trajectories are shown in Appendix -B and supplementary videos.
@@ IV-C1 Plain push-T
Figure 6: World model prediction in GPC for push-T. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 6 shows the quality of world model prediction of push T for GPC-RANK and GPC-OPT. The images shown are all predicted by the world model. They not only have very good visual effect and look realistic, but also they can accurately predict the dynamics reaction from given actions. From the last column, we can see even when the differences between action sequences are very small, the world model can still capture the details and predict the reaction in a meticulous way. This is very important to make GPC work.
IV-C2 Push-T under collision
Next, we hope to investigate how GPC works under more complex scenarios, involving harder tasks and more complex dynamics. We put an object of shape “A” on top of the target position so that when pushing T, it will inevitably collide with object A. The training data is $100$ push-T expert demos with collision of A and some random exploration demos.
Performance. We test $5$ seeds in total, and baseline model, GPC-RANK, GPC-OPT succeed for $2$, $4$, $3$ seeds respectively.
We then proceed to test in an even harder situation: we put objects of shape “A” and “R” on top of the target position at the same time. Case 1: When pushing T, T may collide with A and R at the same time. Case 2: When pushing T, T would collide with one letter (e.g., “R”) and this letter will collide with another letter (e.g., “A”). We train the policy and world model here using $100$ plain push-T expert demonstrations, $100$ push-T colliding A expert demonstrations, and $100$ push-T colliding A&R expert demonstrations.
Performance. In the total of $5$ testing seeds, baseline model, GPC-RANK, GPC-OPT succeed $2$, $3$, $4$ respectively.
Fig. 10(b)(c) show the trajectories for both of these tests. More testing trajectories are shown in Appendix -B and supplementary videos.
@@ IV-C2 Push-T under collision
Figure 7: World model prediction in GPC for push-T with collision of A. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
@@
Figure 8: World model prediction in GPC for push-T with collision of A and R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 7 and Fig. 8 show the quality of world model prediction for GPC-RANK and GPC-OPT, having collisions during push-T. The images shown are all predicted by the world model. With adding some collision objects, the world model is not only able to give near perfect prediction (i.e. no deformation) for a complex scene composing several objects, but also it can accurately predict the dynamics among these objects.
IV-C3 Generalization to unseen cases
Last but not least, we want to understand whether GPC can help when generalizing to unseen cases. We use the policy and world model trained with push-T, push-T colliding A, push-T colliding A&R expert demos. But we perform the test to push T when an object of letter “R” is on top of the target. This situation has never been seen in any of the training data. We test three different starting positions for T and R.
Performance. GPC-RANK and GPC-OPT succeed in all of the three tests while baseline model fails in one of the cases. The reason behind this is that when generalizing to unseen cases, the policy’s performance degrades. So we need to either generate more samples from the action policy’s distribution (like giving more opportunities to generate an optimal proposal) or do gradient ascent to optimize the action proposal.
Fig. 11 shows one test. More testing trajectories are shown in Appendix -B and supplementary videos.
@@ IV-C3 Generalization to unseen cases
Figure 9: World model prediction in GPC for push-T with collision of R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 9 shows the quality of world model prediction of push T under collision in unseen scenarios for GPC-RANK and GPC-OPT. A scenario only composes of T and R is never seen in the training data. However, our world model can still give good predictions for action sequences here. It verifies the ability of generalization for our world model.
Remark 3 (Hallucination).
One of the important issues of generative modeling is hallucination, and our visual world model is no exception. In particular, there are several cases where the world model predicts images of the future that violates rigid-body physics. For example, the objects T, A, and R may penetrate each other, as we show in Appendix -C. Interestingly, even though the predictions can violate physics, using such predictions for planning still leads to GPC’s superior performance. We believe learning a visual world model that strictly satisfies the laws of physics remains a challenging task, and strategies such as data scaling and baking in physics priors may help solve this challenge.
(a) Test for push-T
(b) Test for push-T with collision of A
(c) Test for push-T with collision of A and R
Figure 10: Real-world tests for plain push-T and push-T with collision. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). The first and last column show the initial and final position of the object, respectively. The middle five columns show overlaid trajectories generated by the policies. Results for other real-world tests are shown in Appendix -B and supplementary videos.
Figure 11: Real-world test for out-of-domain scene. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). This displays GPC would help to make policy succeed in unseen scene while baseline failed. Remaining results for push-T with collision of R are shown in Appendix -B and supplementary videos.
V Related Work
@@
VI Conclusion
We presented generative predictive control (GPC), a learning control paradigm that combines a generative policy with a predictive world model during online planning. It is worth noting that GPC bears great similarity with the classical literature on predictive control in the sense that the dynamics model is first identified and then optimization is performed online to decide actions. The new insights offered by GPC, perhaps, are that (a) modern conditional diffusion models provide a powerful tool to learn controllable action-conditioned visual world models; and (b) a generative policy trained from behavior cloning provides good warmstarts for online optimization that seem promising to partially tackle the nonconvexity of online optimization. Through an in-depth investigation on planar pushing—both simulated and real—we demonstrated the superior robustness of GPC in both state-based and vision-based robotic manipulation.
Limitations. Two issues are worth highlighting. The first issue is already mentioned in Remark 3 and concerns the problem of hallucinating future predictions that violate the laws of physics. To address this, it may be possible to leverage the so-called world foundation models [2]—trained on Internet-scale videos—and finetune them for specific robotic tasks. The second issue is associated with the computational complexity of using an expensive diffusion-based world model in online planning. It is worth noting that, although GPC dominates the baseline in terms of success rates, it is generally slower than the baseline due to predicting the future using the world model. A promising solution is to leverage consistency models [51] and one-step diffusion [61] for faster world model prediction.
Acknowledgments
This work was partially supported by the Harvard University Dean’s Competitive Fund for Promising Scholarship. The authors thank William Zhang for his help on hardware design and prototyping. We thank members of the Harvard Computational Robotics Group for their contributions in collecting demonstrations. Their collective support has been instrumental in the completion of this work.

### v2

@@ Appendix D Vision-based Simulation Tasks
Policy. We train a vision-based diffusion policy ${\cal P}(\cdot)$ [7] using ResNet18 as the vision encoder and UNet as the action diffusion network. We train the policy for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon of 2, prediction horizon of 16, and action horizon of 9.
@@
World model. As in §3, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. Same as in §4.1, the training data includes random explorations.
@@
Reward. Different from the state-based planar pushing, since the future predictions are images, we cannot directly compute a reward signal. Therefore, for push-T, we first train a pose prediction network with ResNet18 plus MLP, and then compute the registration loss using the predicted object pose. For triangle drawing, we obtain the reward for expert demonstrations, defined as the number of pixels drawn on the pattern. For block stacking, we obtain the reward for expert demonstrations defined as the distance between two blocks. Then we train a reward predictor to predict the reward from any input images. Note that this learned reward model ${\cal R}(\cdot)$ is still differentiable. Besides, we propose VLM can be used to decide which action proposal to pick in a zero-shot way. We utilize ChatGPT to pick the best from predicted consequences for $10$ actions proposals.
@@
Online planning: With the vision-based policy, the visual world model, and the learned reward model, we follow the same online planning protocol described in §C.
Appendix E Offline RL
It is worth noting that the composition of the reward model ${\cal R}(\cdot)$ and the world model ${\cal W}(\cdot)$ in GPC can be identified as the so-called Q-value function in offline RL, as they predict the reward of an action chunk $a_{t:t+T}$ given the current information $I_{t}$. However, there are three subtle differences between GPC and offline RL. (a) GPC learns an explicit dynamics model instead of just approximating the Q-value function (this is similar to model-based offline RL). (b) The composition of ${\cal R}(\cdot)$ and ${\cal W}(\cdot)$ provides a Q-value function for “action chunks” while the usual Q-value function involves a single action. (c) GPC learns the dynamics model using diverse trajectories (without rewards), instead of only expert demonstrations. In §4, we compare GPC with an offline RL method adapted from single action to action chunk and show that GPC outperforms offline RL.
In experiment section, we compare with an offline RL based planning baseline [31]. While we rely on the world model to rank the action proposals, offline reinforcement learning ranks the action proposals using the value function [31]. It firstly learns the value function via offline RL pre-training, and then performs test-time action re-ranking via the learned value function. To train a value function $Q(I_{t},a_{t:t+T})$ where $I_{t}$ is the state and $a_{t:t+T}$ is the action chunk, we need to obtain a reward function that can be used to supervise the value function training. Consistent with the GPC-RANK method, we use registration error between the object and the target at $I_{t+T+1}$ as the reward. We use the same length of action chunk ($T$) as GPC-RANK for a fair comparison.
Implicit Q-learning (IQL) [41] is used as the offline RL method to pretrain a value function $Q(I_{t},a_{t:t+T})$. For vision-based tasks, we use the pretrained vision encoder (ResNet18) from the action policy to encode the observation images $I_{t}$ to latent vectors $s_{t}$. We then train the value function $Q(s_{t},a_{t:t+T})$. With the trained value function, we can rank and select the action proposal with the highest Q-value.
$a_{t:t+T}^{(k\star)}=\operatorname*{arg\,max}_{a_{t:t+T}^{(k)},k=1\ldots K}Q(s_{t},a_{t:t+T}^{(k)}),$ (A3)
which is the same as the GPC-RANK method in (2).
Appendix F Real-World Experimental Results
Here we provide more results for real-world experiments.
F.1 Test Trajectories
In this section, we provide examples of the test trajectories for five real-world tasks. The test trajectories for push-T with collision of A and R, clothes folding can be found in §4.3. Here we present trajectories for plain push-T, push-T with collision of A, and push-T in an unseen testing scenarios. More videos for all test seeds can be found in https://computationalrobotics.seas.harvard.edu/GPC/.
(a) Test for push-T
(b) Test for push-T with collision of A
Figure A1: Real-world tests for plain push-T and push-T with collision. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). The first and last column show the initial and final position of the object, respectively. The middle five columns show overlaid trajectories generated by the policies.
Figure A2: Real-world test for out-of-domain scene. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). This displays GPC would help to make policy succeed in unseen scene while baseline failed.
F.2 Quality of world model predictions
Fig A3, Fig A4, Fig A5, Fig A6, and Fig A7 show the quality of world model predictions in GPC.
@@ F.2 Quality of world model predictions
Figure A3: World model prediction in GPC for push-T. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
@@
Figure A4: World model prediction in GPC for push-T with collision of A. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
@@
Figure A5: World model prediction in GPC for push-T with collision of A and R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
@@
Figure A6: World model prediction in GPC for push-T with collision of R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Figure A7: World model prediction in GPC for clothes folding. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. The one in red box is picked by policy. To note that only in the selected action proposal, the gripper is grasping the clothes according to the predictions (last image in row 4), while in row 1-3, the gripper, in a closed status, does not managed to grasp anything and in row 5-6, the gripper is still open and not grasping.
Appendix G Related Work